# SynFER: Towards Boosting Facial Expression Recognition with Synthetic Data

### v2

@@ 4.3 Semantic Guidance for Precise Expression Control
Due to the imbalanced distribution of FER labels in the training data and the potential ambiguity between certain facial expressions Zhang et al. (2024b), such as disgust, relying solely on textual and FAU conditions might not guarantee the faithful generation of these expressions. To address this issue, we propose incorporating semantic guidance on the textual embeddings $c^{\text{text}}$, during the later stages of the denoising process. We leverage external knowledge from open-source FER models to steer the generation process, ensuring a more accurate and faithful synthesis of hard-to-distinguish facial expressions.
Layout Initialization. During inference, we select a random face image $x^{s}$ from FEText and invert it to initialize the noise sample $x_{T}^{s}$(Eq. 1). Since early diffusion stages shape the global layout of the image Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023), this strategy helps preserve the natural facial structure, ensuring the generated images are coherent, high-quality, and visually consistent with real-world expressions.
Semantic Guidance. In the early steps of the diffusion process, the generation process is conditioned on the original textual condition $c^{\text{text}}$. To further induce the generation of facial expression images corresponding to their FER labels $y$, we iteratively update the textual condition in the subsequent time steps. Specifically, a facial expression classifier $f(\cdot)$ is utilized for the injection of complex semantics.
To guide the generated images towards the specific class $y$, we propose to do so by updating the textual embeddings. Given an intermediate denoised sample $x_{t}$ at timestep $t$, following Eq. 15 in DDPM Ho et al. (2020), we first estimate the one-step prediction of the original image $\hat{x}_{0}$ as:
@@
$c^{\text{text}}_{t-1}=c^{\text{text}}_{t}+\lambda_{g}\frac{\nabla_{c^{\text{text}}_{t}}\mathcal{L}_{g}}{||\nabla_{c^{\text{text}}_{t}}\mathcal{L}_{g}||_{2}}$ (6)
where $\lambda_{g}$ and $c_{t-1}^{\text{text}}$ denote the step size and the updated textual embedding at timestep $t-1$, respectively. In the latter steps of the diffusion process, the noise estimator network $\epsilon_{\theta}$ is conditioned on the updated textual embeddings rather than the original one.
@@ 4.4 Diffusion-based Label Calibrator (FERAnno)
Image Inversion. To extract facial features and cross-attention maps with the diffusion model $\epsilon_{\theta}$, we first inverse the generated image $x_{0}$ back to the noise sample $x_{t}$ at a denoising timestep $t$, following a predefined scheduler, as described in Eq. 1. To preserve facial details, we set $t=1$ during the inversion process, ensuring that the facial features remain as close as possible to the original generated image $x_{0}$. This partially denoised sample is then passed through the trained denoising network, allowing us to extract rich facial features and cross-attention maps from intermediate layers, which are critical for capturing detailed facial patterns.
Feature Extraction. Given the inverted noise sample $x_{1}$ and the corresponding textual condition $c^{\text{text}}$ and AU condition $c^{\text{au}}$, we can extract the multi-scale feature representations and textual cross-attention maps from the U-Net $\epsilon_{\theta}$ as $\left\{\mathcal{F},\mathcal{A}\right\}=\epsilon_{\theta}(x_{1},t_{1},c^{\text{text}},c^{\text{au}})$, where $\mathcal{F}$ and $\mathcal{A}$ denote the multi-scale feature representations and the cross-attention maps, respectively. $\mathcal{F}$ contains multi-scale feature maps from different layers of the U-Net $\epsilon_{\theta}$ with four different resolutions. $\mathcal{A}$ contains the cross-attention maps drawn from the 16 cross-attention blocks in $\epsilon_{\theta}$. Both the feature representation $\mathcal{F}$ and the cross-attention maps $\mathcal{A}$ are regrouped according to their resolutions.
Multi-scale Features and Attention Maps Fusion. Given that the multi-scale feature maps $\mathcal{F}$ capture global information essential for image generation, and the cross-attention maps provide class-discriminative information as well as relationships between object locations Tang et al. (2022); Caron et al. (2021), FERAnno fuses both features and attention maps within a dual-branch encoder architecture for pseudo-label annotation. An overview of this architecture is shown in Fig. 4.
We first compute the mean of the regrouped attention maps, denoted as $\mathcal{A}_{\text{reg}}$, yielding a set of averaged attention maps $\bar{\mathcal{A}}$. Both the feature maps $\mathcal{F}$ and the averaged attention maps $\bar{\mathcal{A}}$ are then passed through a residual convolution block to prepare them for further processing. To effectively integrate information at different scales, we introduce a bi-directional cross-attention block to fuse the features and attention maps. 1$\times$1 convolutions are employed at various stages to adapt the fusion across multiple resolution layers. Finally, the fused feature maps and attention maps are concatenated and passed through a linear layer, which outputs a probability vector for predicting facial expression classes.
@@ 5 Experiments
We conduct extensive experiments to evaluate both the generation quality of our synthetic data (Sec. 5.1) and its effectiveness in FER tasks (Sec. 5.2). For more details on experimental setup, implementation details are provided in the appendix.
@@ 5.1 Generation Quality
FID ($\downarrow$) HPSv2($\uparrow$) FS($\uparrow$) MPS ($\uparrow$) FER Acc.($\uparrow$) FAU Acc.($\uparrow$) EA ($\uparrow$) FF ($\uparrow$)
@@
SynFER 16.32 0.280 4.26 - 55.14 93.31 59.64 76.79
Table 1: Ours vs.’ shows the proportion of users who prefer our method over the alternative. An MPS above 1.00 and results above 50% in the user study indicate our method outplays the counterpart. FS, FER Acc., FAU Acc., EA and FF denote FaceScore Liao et al. (2024), FER accuracy, facial action unit accuracy, expression alignment and face fidelity, respectively.
We present both objective metrics and subjective user studies, comparing our method to existing state-of-the-art (SOTA) diffusion models Rombach et al. (2022a); Chen et al. (2023a); Li et al. (2024) and the latest facial expression generation technique, FineFace Varanka et al. (2024).
We compute FID between the synthesis images and the test set of the AffectNet Mollahosseini et al. (2017). HPSv2 Wu et al. (2023b) and MPS Zhang et al. (2024a) evaluate the human preferences of the overall synthesis images, while FaceScore Liao et al. (2024) measures the quality of the generated faces. Tab. 1 shows that our method outperforms popular diffusion models and the SOTA facial expression generation method FineFace Varanka et al. (2024), across all metrics of image quality, human preference and facial expression accuracy. Notably, the advantages of SynFER in both FE Acc. and AU Acc. indicate its outstanding controllability in facial expression generation.
@@ 5.2 Effectiveness of Synthetic Dataset
Self-supervised Representation Learning. We trained self-supervised learning (SSL) models, including BYOL Grill et al. (2020), MoCo v3 Chen et al. (2021), and SimCLR Chen et al. (2020), using real-world data, our synthetic data, and a combination of both. The linear probe performance of these models was evaluated on three widely used facial expression recognition (FER) datasets: RAF-DB Li et al. (2017), AffectNet Mollahosseini et al. (2017), and SFEW Dhall et al. (2011), with results reported in Tab. 3. All SSL models were trained with a ResNet-50 architecture He et al. (2016). Notably, state-of-the-art methods in self-supervised facial representation learning, such as MCF Wang et al. (2023), FRA Gao & Patras (2024), and PCL Liu et al. (2023c), were pre-trained on much larger face datasets like LAION-Face Zheng et al. (2022), VGGFace2 Cao et al. (2018), and VoxCeleb Nagrani et al. (2020). However, these models underperformed on FER tasks compared to ours, highlighting that existing large-scale face datasets may lack the high-quality and diverse facial expression patterns required for accurate FER. Results demonstrate that combining real-world and synthetic data consistently boosts SSL baselines. Remarkably, even when MoCo v3 was trained solely on our synthetic data, it achieved a 2.12% improvement on RAF-DB, underscoring the effectiveness of our approach in capturing critical facial expression details that are essential for FER.
Supervised Representation Learning.
@@
Table 4: Performance comparisons with SOTA few-shot learning methods on 5-way few-shot FER tasks with a 95% confidence interval. (*) indicates training with both real-world data and our synthesis data.
Method RAF-DB AffectNet FERPlus
UAR WAR UAR WAR UAR WAR
FaRL 24.98 38.53 26.95 26.95 24.27 35.26
FLAVA 14.35 38.69 14.26 14.26 12.50 28.43
CLIP ViT-B/16 38.66 36.16 34.35 34.35 34.33 45.14
Exp-CLIP ViT-B/16 48.96 54.50 39.98 39.98 40.81 53.02
Exp-CLIP* ViT-B/16 51.23 56.32 41.68 41.68 42.33 56.02
CLIP ViT-L/14 47.22 41.13 34.46 34.47 33.82 46.67
Exp-CLIP ViT-L/14 58.70 65.37 44.27 44.27 48.28 55.42
Exp-CLIP* ViT-L/14 60.41 68.34 46.85 46.85 50.26 57.98
Table 5: Zero-shot performance comparison of CLIP for FER, reporting both Weighted Average Recall (WAR) and Unweighted Average Recall (UAR) as in previous works Zhao et al. (2024). (*) indicates models trained with both real-world and synthetic data.
We validate the effectiveness of SynFER for supervised representation learning by evaluating its performance on RAF-DB and AffectNet (Tab. 3). We compare with SOTA FER models, including Ada-DF Liu et al. (2023a), POSTER++ Mao et al. (2024), and APViT Xue et al. (2022). The results demonstrate that incorporating synthetic data consistently enhances both baseline models and the latest SOTAs in supervised facial expression recognition. Notably, APViT benefits from the synthetic data with improvements of 0.27% on RAF-DB and 0.32% on AffectNet. While the improvements in supervised learning are more modest compared to self-supervised learning, they remain consistent. This is likely due to the stricter distribution alignment required in supervised learning between synthetic training data and real-world test data. In the following section on scaling behavior analysis, we provide further insights, showcasing the use of the distribution alignment technique, Real-Fake Yuan et al. (2023), to alleviate this problem.
Few-shot Learning. Addressing the challenge of limited labeled FER data across different scenarios, we explore the potential of synthetic data to enhance few-shot learning, as presented in Tab. 5. Following the protocol established by CDNet Zou et al. (2022), we train models on five basic expression datasets and evaluate them on three compound expression datasets: CFEE_C Du et al. (2014), EmotionNet_C Fabian Benitez-Quiroz et al. (2016), and RAF_C Li et al. (2017). To benchmark our approach, we compare it against SOTA few-shot learning methods, including InfoPatch Liu et al. (2021), LR+DC Yang et al. (2021), and STARTUP Phoo & Hariharan (2021). The results clearly demonstrate that integrating synthetic data consistently enhances few-shot FER performance across key metrics. This highlights the ability of synthetic data, with its broader range of FER patterns, to bridge the gap in data-limited scenarios, allowing models to better generalize to complex, real-world expressions in few-shot tasks.
Multi-modal Fine-tuning. The synthesis data encompasses multiple modalities, including generated images, textual prompts, and FER labels. To assess its impact on multi-modal fine-tuning for FER, we focus on fine-tuning the vision-language foundation model CLIP Wang et al. (2022), as its performance on face-related tasks is widely regarded as sub-optimal Guo et al. (2023); Chen et al. (2024). Building on Exp-CLIP Zhao et al. (2024), we fine-tune the models on CAER-S Lee et al. (2019) and evaluate their zero-shot performances on the datasets outlined in Tab. 5. Our results show that the inclusion of detailed textual prompts and a larger training image set significantly enhances the generalization ability of Exp-CLIP in understanding facial expressions, achieving significant improvements such as +2.58% UAR on the AffectNet dataset.
@@ 5.3 Ablation Study
Effectiveness of FAU Control. We validate the effectiveness of SynFER by examining how Facial Action Units (FAUs) enhance the generation process and refine facial details. As illustrated in Fig. 5, samples generated with FAU control (third column) exhibit facial expressions that more accurately match their assigned labels compared to those generated with only text guidance (second column). For example, the ’fear’ expression, driven by FAUs like Inner Brow Raiser and Lip Stretcher, becomes more distinct (third column, second row), making it easier to differentiate from other emotions such as ’surprise.’ Similarly, ’disgust’ is more pronounced with FAUs like Lid Tightener. Without FAU control, facial expressions (second column) tend to blur, as different categories show overlapping features. Quantitative results in Tab. 6 highlight the impact of FAU control: FER accuracy increases from 34.62% to 48.74%, and FAU detection accuracy rises from 88.91% to 92.37%. This also translates into improved downstream performance on RAF-DB and AffectNet.
Figure 5: Generated samples. The first and second rows are fear and disgust, respectively.
@@
Table 6: Ablation study on the influence of AU injection and semantic guidance (SG) on both the generation quality and supervised representation learning. SD denotes Stable Diffusion , which is used as a baseline.
Effectiveness of Semantic Guidance. We further explore the impact of semantic guidance (SG) on both generation quality and supervised representation learning, as shown in Fig. 5 and Tab. 6. By updating text embeddings to better align with the target facial expression category, SG improves the accuracy of the generated expressions by 6.4%, compared to static text and FAUs. The samples in the last column of Fig. 5 show more exaggerated facial expressions than those in the third column, with SG enhancing the intensity.
Reliability of FERAnno. We assess the reliability of FERAnno as a label calibrator by evaluating its performance on two FER datasets and visualizing its attention maps in Tab. 3 and Fig. 6. FERAnno significantly outperforms previous SOTAs, achieving a +0.51% improvement on RAF-DB and a +1.34% improvement on AffectNet over the second-best models. The attention maps in Fig. 6 further demonstrate FERAnno’s ability to accurately locate facial expression-related facial features, such as jaw-dropping and furrowed eyebrows, highlighting the diffusion model’s great semantic understanding and fine-grained facial expression recognition.
@@
Synthetic Data Scaling Analysis. Following Tian et al. (2024); Fan et al. (2024), we investigate the scaling behavior of synthetic data in both self-supervised and supervised learning paradigms. To highlight the potential of synthetic FER data, we train models exclusively on synthetic images, without combining real-world data. The results in Fig. 7 (a)-(b) show a stronger scaling effect in self-supervised learning compared to supervised learning, where performance improves significantly with more data. This difference is likely due to the need for better distribution alignment in supervised learning Yuan et al. (2023). While SynFER focuses on addressing FER data scarcity, aligning the synthetic data distribution with real-world data is crucial for supervised tasks. To further explore this, we apply the Real-Fake technique Yuan et al. (2023) for real and synthetic data distribution alignment, and present the results in Fig. 7 (c). Compared to standard supervised learning, Real-Fake demonstrates a clear performance boost.
Figure 7: Scaling up the synthetic FER dataset. MoCo v3 (ResNet-50) Chen et al. (2021) is used for SSL pre-training, and linear probe performance is evaluated on AffectNet and RAF-DB. The SOTA FER model, POSTER++ Mao et al. (2024), is trained using supervised learning (with and without the Real-Fake technique Yuan et al. (2023)) on our synthetic dataset and evaluated on the same two target FER datasets. ★ is model’s performance trained on corresponding real data.

### v3

@@ 4.3 Semantic Guidance for Precise Expression Control
Due to the imbalanced distribution of FER labels in the training data and the potential ambiguity between certain facial expressions [88], such as disgust, relying solely on textual and FAU conditions might not guarantee the faithful generation of these expressions. To address this issue, we propose incorporating semantic guidance on the textual embeddings $c^{\text{text}}$, during the later stages of the denoising process. We leverage external knowledge from open-source FER models to steer the generation process, ensuring a more accurate and faithful synthesis of hard-to-distinguish facial expressions.
Layout Initialization: During inference, we select a random face image $x^{s}$ from FEText and invert it to initialize the noise sample $x_{T}^{s}$(Eq. 1). Since early diffusion stages shape the global layout of the image [87, 61, 55], this strategy helps preserve the natural facial structure, ensuring the generated images are coherent, high-quality, and visually consistent with real-world expressions.
Semantic Guidance: In the early steps of the diffusion process, the generation process is conditioned on the original textual condition $c^{\text{text}}$. To further induce the generation of facial expression images corresponding to their FER labels $y$, we iteratively update the textual condition in the subsequent time steps. Specifically, a facial expression classifier $f(\cdot)$ is utilized for the injection of complex semantics. To guide the generated images towards the specific class $y$, we propose to do so by updating the textual embeddings. Given an intermediate denoised sample $x_{t}$ at timestep $t$, following Eq. 15 in DDPM [32], we first estimate the one-step prediction of the original image $\hat{x}_{0}$ as:
@@
[…]
where $\lambda_{g}$ and $c_{t-1}^{\text{text}}$ denote the step size and the updated textual embedding at timestep t1t-1, respectively. In the latter steps of the diffusion process, the noise estimator network $\epsilon_{\theta}$ is conditioned on the updated textual embeddings rather than the original one.
@@ 4.4 Diffusion-based Label Calibrator (FERAnno)
Image Inversion: To extract facial features and cross-attention maps with the diffusion model $\epsilon_{\theta}$, we first inverse the generated image $x_{0}$ back to the noise sample $x_{t}$ at a denoising timestep $t$, following a predefined scheduler, as described in Eq. 1. To preserve facial details, we set $t=1$ during the inversion process, ensuring that the facial features remain as close as possible to the original generated image $x_{0}$. This partially denoised sample is then passed through the trained denoising network, allowing us to extract rich facial features and cross-attention maps from intermediate layers, which are critical for capturing detailed facial patterns.
Feature Extraction: Given the inverted noise sample $x_{1}$ and the corresponding textual condition $c^{\text{text}}$ and AU condition $c^{\text{au}}$, we can extract the multi-scale feature representations and textual cross-attention maps from the U-Net $\epsilon_{\theta}$ as $\left\{\mathcal{F},\mathcal{A}\right\}=\epsilon_{\theta}(x_{1},t_{1},c^{\text{text}},c^{\text{au}})$, where $\mathcal{F}$ and $\mathcal{A}$ denote the multi-scale feature representations and the cross-attention maps, respectively. $\mathcal{F}$ contains multi-scale feature maps from different layers of the U-Net $\epsilon_{\theta}$ with four different resolutions. $\mathcal{A}$ contains the cross-attention maps drawn from the 16 cross-attention blocks in $\epsilon_{\theta}$. Both the feature representation $\mathcal{F}$ and the cross-attention maps $\mathcal{A}$ are regrouped according to their resolutions.
Multi-scale Features and Attention Maps Fusion: Given that the multi-scale feature maps $\mathcal{F}$ capture global information essential for image generation, and the cross-attention maps provide class-discriminative information as well as relationships between object locations [70, 9], FERAnno fuses both features and attention maps within a dual-branch encoder architecture for pseudo-label annotation. An overview of this architecture is shown in Fig. 4. We first compute the mean of the regrouped attention maps, denoted as $\mathcal{A}_{\text{reg}}$, yielding a set of averaged attention maps $\bar{\mathcal{A}}$. Both the feature maps $\mathcal{F}$ and the averaged attention maps $\bar{\mathcal{A}}$ are then passed through a residual convolution block to prepare them for further processing. To effectively integrate information at different scales, we introduce a bi-directional cross-attention block to fuse the features and attention maps. $\times1$ convolutions are employed at various stages to adapt the fusion across multiple resolution layers. Finally, the fused feature maps and attention maps are concatenated and passed through a linear layer, which outputs a probability vector for predicting facial expression classes.
@@ 5 Experiments
We conduct extensive experiments to evaluate both the generation quality of our synthetic data (Sec. 5.1) and its effectiveness in FER tasks (Sec. 5.2). Implementation details are in the appendix.
Evaluation metrics: In the experimental evaluation, we employ both objective metrics and user studies to comprehensively assess the generation quality. For synthetic image quality, we utilize Fréchet Inception Distance (FID) [82] to measure distribution similarity with real data, FaceScore (FS) [47] for facial quality assessment, and Human Preference Score v2 (HPSv2) [76] with Multi-dimensional Preference Score (MPS) [85] for human perception evaluation. Facial expression accuracy is quantified using pre-trained classifiers and Facial Action Unit Accuracy (FAU Acc.) via AU detection models. For subjective user study, the subjects are asked to select the images with better expression alignment and face fidelity for pairs of synthetic images generated by SynFER and other baselines. Downstream task performance is evaluated through linear probing accuracy in self-supervised learning, classification accuracy in supervised settings, and both Weighted Average Recall (WAR) and Unweighted Average Recall (UAR) for zero-shot recognition. Few-shot learning capabilities are measured using standard n-way k-shot protocols across compound expression datasets.
@@ 5.1 Generation Quality
  FID ($\downarrow$)   HPSv2($\uparrow$)   FS($\uparrow$)   MPS ($\downarrow$)   FER Acc.($\uparrow$)   FAU Acc.($\uparrow$)   EA ($\uparrow$)   $F$ ($\uparrow$)
@@
  SynFER   16.32   0.280   4.26   0.50   55.14   93.31   59.64   76.79
Table 1: Generation quality comparisons. Ours vs.’ shows the proportion of users who prefer our method over the alternative. An MPS above 1.00 and results above 50% in the user study indicate SynFER outplays the counterpart. FS, FER Acc., FAU Acc., EA and $F$ denote FaceScore [47], FER accuracy, facial action unit accuracy, expression alignment and face fidelity, respectively.
We present both objective metrics and subjective user studies, comparing our method to state-of-the-art (SOTA) diffusion models [64, 10, 40] and the latest facial expression generation technique, FineFace [72]. We compute FID between the synthesis images and the test set of the AffectNet [58]. Tab. 1 shows that our method outperforms popular diffusion models and SOTA facial expression generation method FineFace [72], across all metrics of image quality, human preference and facial expression accuracy. Notably, the advantages of SynFER in both FE Acc. and AU Acc. indicate its outstanding controllability in facial expression generation.
@@ 5.2 Effectiveness of Synthetic Dataset
Self-supervised Representation Learning: We trained self-supervised learning (SSL) models, including BYOL [29], MoCo v3 [13], and SimCLR [12], using real-world data, our synthetic data, and a combination of both. The linear probe performances are evaluated on three widely used facial expression recognition (FER) datasets: RAF-DB [43], AffectNet [58], and SFEW [18], with results reported in Tab. 2. All SSL models are trained with a ResNet-50 architecture [31]. Notably, SOTA self-supervised facial representation learning, such as MCF [73], FRA [27], and PCL [51], are pre-trained on much larger face datasets like LAION-Face [89], VGGFace2 [8], and VoxCeleb [59]. However, these models underperformed on FER tasks compared to ours, highlighting that existing large-scale face datasets may lack the high-quality and diverse facial expression patterns required for accurate FER. Results show that combining real-world and synthetic data consistently boosts SSL baselines. Remarkably, even when MoCo v3 was trained solely on our synthetic data, it achieved a 2.12% improvement on RAF-DB, underscoring the effectiveness of our approach in capturing critical facial expression details that are essential for FER.
Supervised Representation Learning: We validate the effectiveness of SynFER for supervised representation learning by evaluating its performance on RAF-DB and AffectNet (Tab. 3). We compare with SOTA FER models, including Ada-DF [50], POSTER++ [56], and APViT [77]. The results demonstrate that incorporating synthetic data consistently enhances both baseline models and the latest SOTAs in supervised facial expression recognition. Notably, APViT benefits from the synthetic data with improvements of 0.27% on RAF-DB and 0.32% on AffectNet. While the improvements in supervised learning are more modest compared to self-supervised learning, they remain consistent. This is likely due to the stricter distribution alignment required in supervised learning between synthetic training data and real-world test data. In the following section on scaling behavior analysis, we provide further insights, showcasing the use of the distribution alignment technique, Real-Fake [83], to alleviate this problem.
@@
Table 4: Comparisons with SOTA few-shot learning methods on 5-way few-shot FER tasks with a 95% confidence interval. (*) indicates training with both real-world and synthetic data.
Few-shot Learning: We explore the potential of synthetic data to enhance few-shot learning, as presented in Tab. 4. Following the protocol from CDNet [92], we train models on five basic expression datasets and evaluate on three compound expression datasets: CFEE_C [20], EmotionNet_C [22], and RAF_C [43]. We compare it against SOTA few-shot learning methods, including InfoPatch [49], LR+DC [79], and STARTUP [62]. The results clearly demonstrate that integrating synthetic data consistently enhances few-shot FER performance across key metrics. This highlights the potential of synthetic data in data-limited scenarios, allowing models to better generalize to complex, real-world expressions in few-shot tasks.
Figure 5: Example of generated samples.
Figure 6: Scaling up the synthetic dataset with MoCo v3 (ResNet-50) [13], and linear probe performance is evaluated on AffectNet and RAF-DB. The SOTA FER model, POSTER++ [56], is trained using supervised learning (with and without the Real-Fake technique [83]) on our synthetic dataset and evaluated on the same two target FER datasets. ★ is model’s performance trained on corresponding real data.
@@ 5.3 Ablation Study
Effectiveness of FAU Control. Fig. 5 shows samples generated with FAU control (third column) exhibit facial expressions that more accurately match their assigned labels compared to those generated with only text guidance (second column). For example, the ’fear’ expression, driven by FAUs like Inner Brow Raiser and Lip Stretcher, becomes more distinct (third column, second row), making it easier to differentiate from other emotions such as ’surprise.’ Similarly, ’disgust’ is more pronounced with FAUs like Lid Tightener. Without FAU control, facial expressions (second column) tend to blur, as different categories show overlapping features. Quantitative results in Tab. 5 highlight the impact of FAU control: FER accuracy increases from 34.62% to 48.74%, and FAU detection accuracy rises from 88.91% to 92.37%. This also translates into improved downstream performance on RAF-DB and AffectNet.
@@
Table 5: Ablation study on AU injection and semantic guidance (SG) on both generation quality and supervised learning. SD denotes Stable Diffusion [65], which is used as a baseline.
Effectiveness of Semantic Guidance: We further explore the impact of semantic guidance (SG) on both generation quality and supervised representation learning ( Fig. 5 and Tab. 5). By updating text embeddings to better align with the target facial expression category, SG improves the accuracy of the generated expressions by 6.4%, compared to static text and FAUs. The samples in the last column of Fig. 5 show more exaggerated facial expressions than those in the third column, with SG enhancing the intensity.
Reliability of FERAnno: We assess the reliability of FERAnno as a label calibrator by evaluating its performance on two FER datasets and visualizing its attention maps in Tab. 3 and Fig. 7. FERAnno outperforms FER SOTAs by +0.51% and +1.34% on RAF-DB and AffectNet over the second-best models. The attention maps (Fig. 7) further show FERAnno’s ability to accurately locate facial expression-related patterns e.g., jaw-dropping and furrowed eyebrows, highlighting its semantic understanding.
@@
Synthetic Data Scaling Analysis: Following [71, 23], we investigate the scaling behavior of synthetic data in both self-supervised and supervised learning paradigms. We train models exclusively on synthetic images, without combining real-world data. The results in Fig. 6 (a)-(b) show a stronger scaling effect in self-supervised learning compared to supervised learning, where performance improves significantly with more data. (c). Compared to standard supervised learning, Real-Fake demonstrates a clear performance boost, which is likely due to the need for better distribution alignment in supervised learning [83]. While SynFER focuses on addressing FER data scarcity, aligning the synthetic data distribution with real-world data is crucial for supervised tasks. To further explore this, we apply the Real-Fake technique [83] for real and synthetic data distribution alignment, and present the results in Fig. 6 (c). Compared to standard supervised learning, Real-Fake demonstrates a clear performance boost.